# Probing Light Bosonic Dark Matter with Transmon Qubits

### v2

@@ III.2 Scalar DM
For the scalar DM $\phi$, the absorption rate can also be derived from the optical theorem by evaluating the self-energies in the medium. Here we adopt the assumption in Ref. Mitridate et al. (2024) that the coupling $g_{\phi}$ is photon-like, and the total absorption rate can be given in the term of the ELF Mitridate et al. (2024),
$R\sim{g_{\phi}^{2}\over e^{2}}{q^{2}\over\omega^{2}}{\rho_{\phi}\over\rho_{T}}{\rm Im}\left[-\frac{1}{\varepsilon(\omega)}\right]$ (10)
@@
We show in the right-panel of the Fig. 1 constraints on the scalar DM in the $g_{\phi}-m_{\phi}$ plane. Again, the region above the contour is excluded. To compare with other constraints from stellar cooling Bottaro et al. (2023) and fifth force Adelberger et al. (2003), one needs to re-parameterize the coupling, $d_{\phi}=g_{\phi}M_{pl}/(\sqrt{4\pi}m_{e})$, where $M_{pl}$ is the Plank mass. Constraints on the new coupling is listed in the left-panel of the Fig. 2.
@@ III.3 Axion-like particle
$R\sim\frac{1}{4m_{a}}\frac{g_{aee}^{2}}{e^{2}}{\omega^{2}\over m_{e}^{2}}\frac{\rho_{a}}{\rho_{T}}{\rm Im}\left[-\frac{1}{\varepsilon(\omega)}\right]$ (11)
where $m_{a}$ is the particle mass of the ALP field, $\rho_{a}$ is the local dark matter density.
We show in the right-panel of the Fig. 2 the exclusion limit on the $g_{aee}$, where the shaded region is excluded. Here we do not consider the axion coming from the sun, and take $\rho_{a}\sim 0.4~{}{\rm GeV/cm^{3}}$ as the local energy density of the ALP on the earth. For the mass range that we care, there are other constraints from solar neutrinos Giudice et al. (2012), red giants Gondolo and Raffelt (2009), as well as solar axion measurement at the XENONnT and the PandaX4T Aprile et al. (2022); Meng et al. (2021), which have $g_{aee}<2.82\times 10^{-11},<1.9\times 10^{-12}$ and $<4.38\times 10^{-12}$, respectively. These constraints are much stronger than that from the current measurement result of the transmon qubit, and we only show the constraint of the electron g-2 in the plot. Although this exclusion limit is not stronger than limits from other observations, it provides the first direct detection limit on the ALP-matter coupling in the meV-eV mass range. Further experiments with more precise measurement and more systematic estimation of background will be able to explore more competitive parameter space.

### v3

@@ III.2 Scalar DM
For the scalar DM $\phi$, the absorption rate can also be derived from the optical theorem by evaluating the self-energies in the medium. Here we adopt the assumption in Ref. Mitridate et al. (2024) that the coupling $d_{\phi ff}$ is photon-like, and the total absorption rate can be given in the term of the ELF Mitridate et al. (2024),
$R\sim{d_{\phi ee}^{2}\over e^{2}}{q^{2}\over\omega^{2}}{\rho_{\phi}\over\rho_{T}}{\rm Im}\left[-\frac{1}{\varepsilon(\omega)}\right]$ (10)
@@
We show in the right-panel of the Fig. 1 constraints on the scalar DM in the $d_{\phi ee}-m_{\phi}$ plane. Again, the region above the contour is excluded. To compare with other constraints from stellar cooling Bottaro et al. (2023) and fifth force Adelberger et al. (2003), one needs to re-parameterize the coupling, $d^{\prime}_{\phi ee}=d_{\phi ee}M_{pl}/(\sqrt{4\pi}m_{e})$, where $M_{pl}$ is the Plank mass. Constraints on the new coupling is listed in the left-panel of the Fig. 2.
@@ III.3 Axion-like particle
$R\sim\frac{1}{4}\frac{g_{aee}^{2}}{e^{2}}{\omega^{2}\over m_{e}^{2}}\frac{\rho_{a}}{\rho_{T}}{\rm Im}\left[-\frac{1}{\varepsilon(\omega)}\right]$ (11)
where $\rho_{a}$ is the local energy density of the ALP.
We show in the right-panel of the Fig. 2 the exclusion limit on the $g_{aee}$, where the shaded region is excluded. Here we do not consider the axion coming from the sun, and take $\rho_{a}\sim 0.4~{\rm GeV/cm^{3}}$ as the local energy density of the ALP on the earth. For the mass range that we care, there are other constraints from solar neutrinos Giudice et al. (2012), red giants Gondolo and Raffelt (2009), as well as solar axion measurement at the XENONnT and the PandaX4T Aprile et al. (2022); Meng et al. (2021), which have $g_{aee}<2.82\times 10^{-11},<1.9\times 10^{-12}$ and $4.38\times 10^{-12}$, respectively. These constraints are much stronger than that from the current measurement result of the transmon qubit, so we only show the constraint of the electron g-2 in the plot, since it has encompassed all the parameter space detectable by the transmon qubit. Although this exclusion limit is not stronger than limits from other observations, it provides the first direct detection limit on the ALP-matter coupling in the meV-eV mass range. Further experiments with more precise measurement and more systematic estimation of background might enable us to explore more competitive parameter space.
Before proceeding to the discussion, Let’s comment on formulas used in our analysis. We acknowledge that the standard Lindhard/Mermin formalism used in DarkELF does not automatically incorporate the complex effects of superconductivity, such as the formation of a gap in the density of states. Our approach provides a conservative, order-of-magnitude estimate for the absorption rate for dark matter. A more precise treatment is beyond the scope of the current version of DarkELF. This simplification is noted as an area for future refinement in more dedicated detector studies. Furthermore, The derivation of the absorption rate is based on some specific assumptions following the strategy of Ref. [40]. It works for current magnitude estimation of constraints, and needs further dedicated calculation when we intend to test a specific ultralight DM model using a transmon qubit device.
For the device performance, there are fundamental and readout related noise sources as noted in Ref. [73], such as telegraph noise, Fano noise, system and amplifier noise, etc. We briefly address mechanisms for mitigating noises. For the electromagnetic noise, one may use high-conductivity materials for Faraday cages to suppress external radio frequency interference, employ coaxial shielding or cryogenic coaxial cables for signal lines to minimize crosstalk and radiation noise, use differential signal transmission to enhance robustness against common-mode electromagnetic interference. For the thermal noise, one may employ dilution refrigerators to cool temperatures to the mK range, suppressing thermally excited quasiparticles, design quasiparticle trap structures to accelerate recombination of thermally generated quasiparticles, use quantum impedance matching to compress thermal noise spectral density to the quantum limit.